OpenStudy (anonymous):

Eleven less than four times the smaller of two consecutive odd integers is three times the larger integer. Find the two integers.

OpenStudy (anonymous):

ill give metals!

OpenStudy (anonymous):

Let x = first odd integer (smaller) x+2 = second odd integer (larger) 3(x+2) = 4x-5 3x+ 6 = 4x -5 Subtract 3x from each side: 3x-3x + 6 = 4x-3x -5 6=x-5 Add +5 to each side: 6+5= x-5+5 11=x x+2= 13.

OpenStudy (anonymous):

Smaller = 11, Larger = 13. Check: Three times the larger is five less than 4 times the smaller. 3(13)= 4(11) -5 39= 44-5

OpenStudy (anonymous):

it says that you answer is wrong :/

OpenStudy (anonymous):

oh sorry i tried my best to help you

OpenStudy (anonymous):

its okay thanks anyways